# thanks: These authors contributed equallythanks: These authors contributed equallythanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ Abstract
Despite extensive investigations, many aspects of charge density waves (CDWs) remain elusive, especially the relative roles of electron-phonon coupling and Fermi surface nesting as the underlying driving mechanisms responsible for the emergence of the CDW vector $\boldsymbol{Q}_{CDW}$. It is puzzling that even though electrons interact strongly with optical phonons in many correlated systems, the actual mode softening is of an acoustic mode. Here we consider monolayer 1H-NbSe2 as an exemplar system, and through an accurate computation of the phonon self-energy, including its off-diagonal components. We provide compelling evidence that the relevant mode is a longitudinal optical phonon that softens by anti-crossing several intervening phonon bands. We also show that $\boldsymbol{Q}_{CDW}$ is fixed by the convolution of the susceptibility and electron-phonon coupling, and that the softened phonons are circularly polarized.
I Introduction
@@
Transition-metal dichalcogenides (TMDCs) are well-known for their layered structures in which transition-metal atoms are sandwiched between two layers of chalcogen, creating a structure that supports a wide variety of electronic interactions driven by valence electron configurations and the spatial symmetries of the unit cell  [TMDC1, TMDC2, TMDC3]. In these compounds, the resulting physical phenomena include the coexistence of charge density wave (CDW) and superconductivity (SC) in NbSe2 [ScCdwExp, ScCdw2]. The simultaneous presence of CDW and SC phases offers a unique platform for exploring effects of complex electron correlations in condensed matter systems.
@@
Among the TMDCs, NbSe2 is distinguished as it exhibits an unusually rich tapestry of properties in both the bulk and monolayer phases. Bulk NbSe2 hosts a CDW with a periodic lattice distortion (PLD) [FengZheng] below $T_{\text{CDW}}\approx 33$ K, along with a coexisting SC phase with $T_{c}\approx 7.2$ K. As the number of layers in a film of NbSe2 is reduced the CDW transition becomes stronger and superconductivity weakens. In the 1H-monolayer limit $T_{\text{CDW}}\approx 145$ K and $T_{c}\approx 1.9$ K [ScCdw3, ScCdw4]. Notably, the layered architecture of TMDCs like NbSe2 also makes them amenable to strain engineering and the construction of heterostructures [strain, hetero1, hetero2]. This adaptability enables a wide range of applications in nanoelectronics, optoelectronics, and quantum devices [Vaskivskyi2015Fast, Akinwande2017Two, Clarke2008Superconducting], where manipulating electronic properties at the nanoscale is of key importance.
@@
The nature of electron-phonon coupling varies across the landscape of materials. In some cases, electron density gets localized and distorts the lattice around the site, effectively enhancing the electronic mass, while in other cases the phonon modulates the bond lengths, strengthening some hoppings and weakening others[polaron1, polaron2]. In polaron physics the former are known as Holstein polarons and the latter as Peierls polarons.[Berciu], with drastic differences in their properties. For Holstein polarons – with increasing electron-phonon coupling, electrons become increasingly more trapped in their polaron clouds and eventually undergo a metal-insulator transition[MIT]. In contrast, for the Peierls polarons, the bonds with strengthened hoppings can link up along certain chains, so that the polaron mass can remain small even for strong electron-phonon coupling. Holstein polarons thus unlikely candidates for bipolaron superconductivity[Mott], but this remains a possibility for Peierls polarons.
This difference should persist in the CDW physics as well. For example, many high-frequency oxygen modes in the cuprates are of Peierls form, modulating copper-to-oxygen hopping, so that the CDW transition would be expected to involve the softening of these optical modes[CuprateSoften1, CuprateSoften2, CuprateSoften3], although this softening has been observed only rarely in experiments.
@@
The softening of an optical phonon to zero energy leads to a form of “quantum strangeness” [Strangeness] in the sense that the soft mode Kohn anomaly[Kohn] must cross all lower-lying phonon modes. Modes of different symmetries can cross without interactions, but those of the same symmetry (irreducible representation (irrep) of the dynamical matrix), interact to open an energy gap or ‘anticross’, where eigenfunctions interchange their characters, see Figure 1(a). At critical softening the Kohn anomaly thus resembles a “Kohn ladder” or a Kohn anomaly which is broken up into many segments separated by anticrossings at ascending energies.
@@
Such a ladder has been predicted[Gutzcharge] and observed[Ding] in the cuprates, where the resulting CDW is consistent with a known optical phonon symmetry of the parent compound[Forgan]. In NbSe2, an experimental study examined but ultimately rejected the possibility of a Kohn ladder[Rez]. However, in monolayer 1H-NbSe2, which lacks inversion symmetry, two possible CDW distortions have been reported, which are consistent with breathing-in and breathing-out distortions of an optical phonon[NbSe2]. Here we re-explore this issue by examining the contribution of electron-phonon coupling to the phonon self-energy to uncover what controls the CDW $q$-vector.
II NbSe2
@@
NbSe2 is a canonical metallic superconducting CDW material, with the superconducting state occurring at a temperature $T_{c}\ll T_{\text{CDW}}$ [Wilson1975, Moncton1977]. Near $T_{\text{CDW}}$, there is a strong Kohn anomaly associated with a soft phonon and a static CDW exhibiting a $3\times 3$ PLD in real space [EPH3, EPH4]. This corresponds to a triple CDW wavevector $\mathbf{Q}_{\text{CDW}}\approx\frac{2}{3}\overline{\Gamma M}$ [Soumyanarayanan2013, EPH5].
@@
Unlike NbS2, NbSe2 does not exhibit magnetic response and it is not close to any magnetic instability in its CDW phase [magnetic, magnet2]. The electronic spin-orbit energy splitting at the Fermi level is negligibly small compared to the CDW energy scale [Johannes2006]. The Fermi surface of NbSe2 is strongly anisotropic, composed of pockets around the $\Gamma$ and $K$ points, preventing perfect nesting at the Fermi energy [Flick, FengZheng, FermiGap]. These features of the electronic structure imply that the PLD-CDW instability is far from conventional being and that it results from a combination of strong electron-phonon coupling and Fermi surface topology [topo1].
@@
Survival of the superconducting state in the presence of such a strong CDW transition indicates that the CDW and superconducting regions avoid each other, confirming a strongly anisotropic Fermi surface within the CDW state [new_nbse2, 2new_nbse2]. This unconventional picture is distinct from the commonly observed single-acoustic-mode softening accompanied by the appearance of a uniform CDW gap around the Fermi surface. Indeed, density functional theory (DFT) based calculations show that both longitudinal acoustic and longitudinal optical phonon modes have strong coupling with electrons on the Fermi surface, and that spin-orbit effects do not play a major role.
@@
However, many questions about the origin of the CDW in NbSe2 remain controversial  [Johannes2006, Borisenko2009, Kiss2007, Calandra2009, Plummer]. In this study, we will focus on the $T_{c}\ll T_{\text{CDW}}\leq T$ regime, and employ DFT-based calculations to address two questions: How does the optical-acoustic phonon mode mixing relate to the Kohn anomaly? And is $\mathbf{Q}_{\text{CDW}}$ determined by the peak in the charge susceptibility or by the anisotropic electron-phonon coupling?
@@
Figure 1: (color online) Phonon dispersions in NbSe2. (a) $\sigma$-dependent phonon spectrum. (b) $\sigma$-dependent self-energy spectrum for acoustical bands. (c) $\sigma$-dependent self-energy spectrum for optical bands. (d) Variation of unit cell parameters vs $\sigma$. Blue line gives the change in the lattice parameter, while red line gives the distance between Nb and Se atoms with respect to the change in $\sigma$.
III Calculating electron-phonon coupling
@@
In conventional many-body approaches, the bare phonon Green’s function is calculated without electron-phonon interactions, which are incorporated subsequently in the phonon self-energy [Mahan]. In contrast, DFT calculations of the phonon spectra do not make an explicit separation into bare and dressed phonons, which complicates comparisons between various phonon results  [Giustino2017, Marini2001, Heid2010, Harkonen2020].
@@
We first use Quantum Espresso (QE) [QE_refs1, QE_refs2] for DFT-based structural optimization and electronic structure computations, see Appendix D for details. A Fermi-Dirac type smearing $\sigma$ is invoked to simulate thermal effects and define an effective temperature $T_{\sigma}=\sigma/k_{B}$, where $k_{B}$ is Boltzmann’s constant. We then use density-functional perturbation theory (DFPT), as implemented in the QE software[QuantEspRenorm], to directly phonon frequencies as a function of $\sigma$. Electron-phonon coupling matrix elements are then computed by using the EPW package [EPW]. The original EPW code gives the absolute value of the electron-phonon coupling and the diagonal elements of the self-energy matrix. To assess the couplings between the phonon modes and the phonon self-energy, we modified the EPW code[EPWmodified] to calculate the off-diagonal elements $\lambda\neq\lambda^{\prime}$ of the phonon self-energy[typy] and the coupled phonon eigenstates.
@@
Figure 2: (color online) Decomposition of self-energy to determine the origin of the $q$-vector close to the instability temperature, $\sigma=0.00565$ Ry. Top row (left to right): (a) Phonon dispersions for various modes, numbered with different colors. (b) Susceptibility, $\chi(q)$ (dashed red line, right-axis)), and $k$-averaged electron-phonon coupling $G(q)$ (remaining curves, left axis). (c) $G(q)\chi(q)$. (d) Diagonal self-energy $\Pi^{\prime}_{\lambda\lambda}$. (e,f) Real (e) and imaginary (f) parts of the off-diagonal self-energy $\Pi^{\prime}_{0\lambda}$, where 0 denotes the soft-mode phonon mode (blue). (g) $1/\Omega$. and (h) $\Pi$. Position of the soft-mode $\bf Q_{CDW}$-vector (solid black vertical lines) and $\bf q=2/3\,\overrightarrow{\Gamma M}$ (dashed black vertical lines) are shown.
Figure 2 presents the phonon spectrum along the $\Gamma-M$ in panel (a), while the remaining frames provide results for various components of the calculations to gain insight into the nature and origin of the CDW instability. Modes 3, 6, and 8 have predominantly longitudinal symmetry and interact strongly with each other. Modes 2 and 7 are predominantly transverse, but since the modes are strictly longitudinal or transverse only along the high-symmetry lines such as $\Gamma-M$, these longitudinal and transverse modes interact away from the high-symmetry lines. These are the only modes with strong electron-phonon coupling, and therefore, we will we focus mainly on these modes. The sharp dip in the phonon dispersion of mode 3 in panel (a) is the Kohn anomaly that drives the CDW instability[Kohn].
@@
Figure 1 summarizes the evolution of the phonon spectra (a-c) and lattice parameters (d) for various $\sigma$ values. Strong $T$-dependence of the phonon frequencies and the appearance of phonon softening are clear evidence that the interactions are inseparably incorporated in the DFT calculations. Strong $T$-dependence of the atomic positions and lattice constants in the numerical results shown in Figure 1(d) suggest the importance of multiphonon processes (noninteracting phonons do not involve thermal lattice expansion). The the phonon self-energy is calculated non-perturbatively here in terms of dressed phonons, which leaves open the questions: How does one extract the bare phonon frequencies? And is the DFT based self-energy consistent with the field theory renormalization of the bare phonon frequencies?
@@
To reconstruct the dynamical matrix in the presence of electron-phonon coupling, we start with a general framework where multiple phonon modes interact with multiple electron bands in the vicinity of the Fermi energy. [Mode and band are used in the Born-Oppenheimer sense where the ionic and electronic degrees of freedom are decoupled.] The electron-phonon interaction is handled perturbatively, see Appendix B for details.
@@
In the multi-mode phonon basis[Giustino1], the renormalized phonons are given by the poles of the interacting Green’s function:
@@
where $\boldsymbol{\omega}_{\boldsymbol{q}}$ is a diagonal matrix with elements $\omega_{\lambda,\boldsymbol{q}}$ in the phonon eigen-mode space describing the bare-phonon frequencies and ${\boldsymbol{\Pi}}(\boldsymbol{q},\omega)$ is the self-consistent electron-phonon self-energy given by,
@@
$G_{\lambda,\lambda^{\prime}}(n,\boldsymbol{k};m,\boldsymbol{k}+\boldsymbol{q})=\mathfrak{g}_{n,m,\lambda}(\boldsymbol{k},\boldsymbol{k}+\boldsymbol{q})\,\mathfrak{g}_{m,n,\lambda^{\prime}}(\boldsymbol{k}+\boldsymbol{q},\boldsymbol{k})$
@@
and $g_{n,m,\lambda}(\boldsymbol{k};\boldsymbol{k}+\boldsymbol{q})$ is the microscopic complex electron-phonon coupling given by[Giustino1]
$g_{n,m,\lambda}(\boldsymbol{k};\boldsymbol{k}+\boldsymbol{q})=\langle|\Psi_{\boldsymbol{k}+\boldsymbol{q},m}|\nabla_{\boldsymbol{q}}V_{\boldsymbol{q}}.\boldsymbol{\xi}_{\lambda\boldsymbol{q}}|\Psi_{\boldsymbol{k},n}\rangle.$ (5)
where $|\Psi_{\boldsymbol{k},n}\rangle$ are the electron Bloch states, see Appendix B for details. Inhomogeneous $\boldsymbol{k}$ dependence in Eq.(5) and hence in Eq.(3) is the result of the anisotropy and the sensitivity of the Bloch wave functions to the crystal structure. In NbSe2 the relevant sections of the Fermi surface responsible for the strong electron interaction are in the K-pockets near the zone boundaries.
Note that $\Pi_{\lambda\lambda^{\prime}}$ depends on the bare phonon frequencies, $\omega_{\lambda\boldsymbol{q}}$, but that an exact solution for the dressed phonons can be obtained in principle by replacing self-consistently the bare terms by the dressed terms [Hedin, Mahan]. Below, we will refer to the replacement of $\omega_{\lambda\boldsymbol{q}}$ by $\Omega_{\lambda\boldsymbol{q}}$ as the Hedin model.
We now discuss how the $\boldsymbol{k}$ dependence of $g_{n,m,\lambda}(\boldsymbol{k};\boldsymbol{k}+\boldsymbol{q})$ drives an unconventional phonon softening and CDW formation distinct from the Peierls mechanism[gruner1994density].
In the limit of negligibly weak lattice potential, the Bloch states can be approximated by plane waves, so that $g_{n,m,\lambda}(\boldsymbol{k};\boldsymbol{k}+\boldsymbol{q})$ in Eq.(5) becomes dependent only on the momentum exchanged with the phonons[Grimvall]. The phonon self-energy in Eq.(2) can then be written as:
@@
The second term on the right hand side of Eq. 6 is closely related to the Lindhard electronic susceptibility (dropping the band index n and considering the spin degeneracy factor 2),
which we calculate using Maximally Localized Wannier Functions (MLWFs) [Wannier]. We further simplify the computation by taking the adiabatic limit ($\omega\to 0$). [Non-adiabatic corrections in NbSe2 are found to be small.] Effects of the $\boldsymbol{k}$ dependence of $g_{n,m,\lambda}(\boldsymbol{k};\boldsymbol{k}+\boldsymbol{q})$ can be seen in $\Pi_{\lambda,\lambda^{\prime}}$ evaluated using the limiting cases given by Eqs. (2) and (6).
Figure 2(a) shows the DFPT phonon dispersions[Baroni] calculated at a temperature slightly above $T_{CDW}$. The CDW transition is accurately predicted at the wavevector $\boldsymbol{Q}_{CDW}\simeq 2\,{\overrightarrow{\Gamma M}}/3$ which agrees with the experimental value, see Appendix A for details. It is convenient to introduce a new quantity here:

### v2

# Giant Kohn anomaly and chiral phonons in the charge density wave phase of 1H-NbSe2: impact of phonon anticrossing

@@ Abstract
Despite extensive investigations, many aspects of charge density waves (CDWs) remain elusive, especially the relative roles of electron-phonon coupling and Fermi surface nesting as the underlying driving mechanisms responsible for the emergence of the CDW vector $\boldsymbol{Q}_{CDW}$. It is puzzling that even though electrons interact strongly with optical phonons in many correlated systems, the actual mode softening is of an acoustic mode. Here we consider monolayer 1H-NbSe2 as an exemplar system, and through an accurate computation of the phonon self-energy, including its off-diagonal components, we provide compelling evidence that the relevant mode is a longitudinal optical phonon that softens by anti-crossing several intervening phonon bands, i.e. a Kohn ladder which has been only observed previously in high temperature superconductors. We also show that $\boldsymbol{Q}_{CDW}$ is fixed by the convolution of the susceptibility and electron-phonon coupling, and that the softened phonons are circularly polarized.
keywords:
NbSe2, chiral, Kohn anomaly, anticrossing, charge-density wave, superconductivity
[…]
In the race to miniaturize – connecting ever smaller semiconductor devices18 with ever thinner metallic wires84 – quantum effects have evolved from being a small perturbation to becoming a driving force for innovation and new devices19, including quantum computers based on quantum entanglement70. Moreover, improved density functional theories (DFTs)45, 85 are advancing our ability to describe exotic band structures such as the linearly dispersing Dirac and Weyl bands37, 4, and providing a firm basis for material-specific many-body models of quantum materials.
Anticrossing of bands is an unusual effect, an example of the strangeness of quantum mechanics, where two bands of the same symmetry cannot cross96, but must maintain a minimum gap between them40. Thus, as the offset parameter between two bands is continuously changed, the bands first move in k-space towards each other, but instead of crossing they appear bounce off of each other and then start moving in the opposite direction with increasing separation. What really happened is that, while the bands were hovering near their minimal separation, they were gradually exchanging their characters, so that far from the minimal gap it looks as if they have simply crossed each other. This is a well-known effect of the off-diagonal matrix elements of the phonon dynamic matrix. An interesting manifestation of anticrossing arises when allowed crossings lead to Dirac-like dispersions and the band continuity can only be demonstrated by passing through more than one Brillouin zone.71 These effects are readily captured for electronic bands in DFT, and, as we shall show, for phononic bands in density functional perturbation theory (DFPT).
We expect anticrossing to also play a key role in one of the most important properties of phonon bands – the driving of phase transitions via soft phonons. The softening is often caused by electron-phonon interaction, leading to a sharp depression of the phonon band in a narrow region of the momentum space, which results in Kohn anomalies50 from the peaks in the electron susceptibility. In many materials, the strongest electron-phonon coupling is induced by optical modes involving ligand ions, which cause large variations of electron hopping parameters. However, when an optical mode goes soft, it produces not a single Kohn anomaly but a stack of anomalies, that we refer to as a Kohn ladder, see LABEL:SM1 and Fig. LABEL:fig:1. The soft phonon must anticross several intermediate bands, breaking the Kohn anomaly into a ladder of several disconnected steps. By extracting the off-diagonal components of the phonon self-energy we can separate the bare phonons at high temperatures from the dressed phonons at lower temperatures, and clearly demonstrate how the displacement pattern and the strong electron-phonon coupling are passed from band to band as the phonon softens. Here, we analyze the monolayer form of the well-known charge-density wave material, NbSe2, by revising the Quantum Espresso (QE) 29, 28 code to expose the underlying phonon anticrossing effects.
@@
Transition-metal dichalcogenides (TMDCs) are well-known for their layered structures in which transition-metal atoms are sandwiched between two layers of chalcogens, creating a structure that supports a wide variety of electronic interactions driven by valence electron configurations and the spatial symmetries of the unit cell  61, 48, 82. The resulting phenomena include the coexistence of charge density wave (CDW) and superconductivity (SC) in NbSe2 15, 56. The simultaneous presence of CDW and SC orders offers a unique platform for exploring effects of complex electron correlations in condensed matter systems.
@@
Among the TMDCs, NbSe2 is distinguished as it exhibits an unusually rich tapestry of properties in both the bulk and thin films. Bulk NbSe2 hosts a CDW with a periodic lattice distortion (PLD) 98 below $T_{\text{CDW}}\approx 33$ K, along with a coexisting SC phase with $T_{c}\approx 7.2$ K. As the number of layers in a film of NbSe2 is reduced, the CDW transition becomes stronger, and superconductivity weakens. In the 1H-monolayer limit $T_{\text{CDW}}\approx 145$ K and $T_{c}\approx 1.9$ K 86, 93. Notably, the layered architecture of TMDCs also makes them amenable to strain engineering and the construction of heterostructures 77, 27, 3. This adaptability enables a wide range of applications in nanoelectronics, optoelectronics, and quantum devices 87, 1, 16, where manipulation of electronic properties at the nanoscale plays a key role.
@@
The nature of electron-phonon coupling varies across the landscape of materials. In some cases, electrons become localized and distort the lattice around the atomic sites, effectively enhancing the electronic mass, while in other cases the phonon modulates the bond lengths, strengthening some hoppings and weakening others97, 81: In polaron physics, the former are known as Holstein polarons and the latter as Peierls polarons95, with drastic differences in their properties. For Holstein polarons, with increasing electron-phonon coupling, electrons become increasingly more trapped in their polaron clouds and eventually undergo a metal-insulator transition68. In contrast, for the Peierls polarons, the bonds with strengthened hoppings can link up along certain chains, so that the polaron mass can remain small even for strong electron-phonon coupling. Holstein polarons are thus unlikely candidates for bipolaron superconductivity2, but this remains a possibility for Peierls polarons. These differences should persist in the CDW physics as well. For example, many high-frequency oxygen modes in the cuprates are of Peierls form, modulating copper-to-oxygen hopping, so that the CDW transition would be expected to involve the softening of these optical modes43, 80, 46, although this softening has been observed only rarely in experiments.
@@
The softening of an optical phonon to zero energy leads to a form of “quantum strangeness” 92 in the sense that the soft mode Kohn anomaly must cross all lower-lying phonon modes. Modes of different symmetries can cross without interactions, but those of the same symmetry (irreducible representation (irrep) of the dynamical matrix), interact to open an energy gap or ‘anticross’, where eigenfunctions interchange their characters. We define a “Kohn ladder” as a Kohn anomaly which is broken up into many segments separated by anticrossings at ascending energies. In LABEL:SM1, we discuss a simple toy model to demonstrate how in such a ladder the phonon displacement pattern is transferred from band to band as the LO phonon softens, so that when a CDW develops, its displacement pattern resembles that of the LO phonon.
@@
Such a Kohn ladder has been predicted63 and observed55 in the cuprates, where the resulting CDW is consistent with the known optical phonon symmetry of the parent compound26. In NbSe2, an experimental study examined but ultimately rejected the possibility of a Kohn ladder88. However, in monolayer 1H-NbSe2, which lacks inversion symmetry, high-pressure calculations report two possible CDW distortions, which are consistent with breathing-in and breathing-out distortions of an optical phonon14. Here we re-explore this issue by examining the contribution of electron-phonon coupling to the phonon self-energy to uncover what controls the CDW $q$-vector.
1 Results/Discussion
1.1 NbSe2
@@
NbSe2 is a canonical metallic superconducting CDW material, with the superconducting state occurring at a temperature $T_{c}\ll T_{\text{CDW}}$ 91, 67. Near $T_{\text{CDW}}$, there is a strong Kohn anomaly associated with a soft phonon and a static CDW exhibiting a $3\times 3$ PLD in real space 89, 41. This corresponds to a triple CDW wavevector $\mathbf{Q}_{\text{CDW}}\approx\frac{2}{3}\overline{\Gamma M}$ 83, 94.
@@
Unlike NbS2, NbSe2 does not exhibit magnetic response and it is not close to any magnetic instability in its CDW phase 53, 99. The electronic spin-orbit splitting at the Fermi level is negligibly small compared to the CDW energy scale 47. The Fermi surface of NbSe2 (Fig. LABEL:fig:deconv(a) Right) is strongly anisotropic, composed of pockets around the $\Gamma$ and $K$ points, preventing perfect nesting at the Fermi energy 25, 98, 11. These features of the electronic structure imply that the PLD-CDW instability is far from conventional and that it results from a combination of strong electron-phonon coupling and Fermi surface topology 44.
@@
Survival of the superconducting state in the presence of such a strong CDW transition indicates that the CDW and superconducting regions in $k$-space avoid each other, confirming a strongly anisotropic Fermi surface within the CDW state 17, 78. This unconventional picture is distinct from the commonly observed single-acoustic-mode softening accompanied by the appearance of a uniform CDW gap around the Fermi surface. Indeed, density functional theory (DFT) based calculations show that both longitudinal acoustic and longitudinal optical phonon modes have strong coupling with electrons on the Fermi surface, and that spin-orbit effects do not play a major role.
Despite decades of study, the microscopic origin of the $\times3$ CDW in NbSe2 remains contested: simple Fermi-surface nesting alone appears insufficient 47; ARPES reveals two gaps and Fermi-surface ‘arcs,’ challenging a uniform-gap, nesting-only picture10; other work highlights strongly momentum-dependent electron–phonon coupling and its interplay with superconductivity49; cross-material analysis supports EPC-driven CDWs as a viable route13; and broader surveys show multiple CDW classes (Peierls, excitonic, EPC-dominated), underscoring why NbSe2 remains debated 100.
@@
In this study, we will focus on the $T_{c}\ll T_{\text{CDW}}\leq T$ regime, and employ DFT-based calculations to address two questions: How does the optical-acoustic phonon mode mixing relate to the Kohn anomaly? And is $\mathbf{Q}_{\text{CDW}}$ determined by the peak in the charge susceptibility or by the anisotropic electron-phonon coupling?
@@
Figure 1: (color online) Phonon dispersions in NbSe2. Phonon branch indices are identified with different colors in the legend. (a) $\sigma$-dependent phonon spectrum. (b) $\sigma$-dependent self-energy spectrum for acoustical bands. (c) $\sigma$-dependent self-energy spectrum for optical bands. (d) Variation of unit cell parameters vs $\sigma$. Blue line gives the change in the lattice parameter, while red line gives the distance between Nb and Se atoms with respect to the change in $\sigma$.
1.2 Calculating electron-phonon coupling
@@
In conventional many-body approaches, the bare phonon Green’s function is calculated without electron-phonon interactions, which are incorporated subsequently in the phonon self-energy 60. In contrast, DFT calculations of the phonon spectra do not make an explicit separation into bare and dressed phonons, which complicates comparisons between various phonon results  30, 62, 39, 36. Here we focus on the DFT calculations, while in LABEL:SM2 we discuss differences with the many-body approach and an approximate way of relating the two approaches.
@@
We use Quantum Espresso (QE) 29, 28 for DFT-based structural optimization and electronic structure computations, see Supporting Information (SI)-1 for details. A Fermi-Dirac type smearing $\sigma$ is invoked to simulate thermal effects and define an effective temperature $T_{\sigma}=\sigma/k_{B}$, where $k_{B}$ is Boltzmann’s constant. We then use density-functional perturbation theory (DFPT), as implemented in the QE software (In the renormalization of phonons by electron-phonon coupling, Quantum Espresso does not consider a full diagonalization in the multimode phonon subspace which would require mode-mode coupling of the phonons.76 Instead QE considers the renormalization at a given phonon mode fixed in the absence of the electron-phonon interaction. In the presence of a number of phonon modes coupled to a single electron branch, the mode-mode coupling can make significant contributions to the phonon softening.), to directly calculate phonon frequencies as a function of $\sigma$. Electron-phonon coupling matrix elements are then computed by using the EPW package 54. The original EPW code gives the absolute value of the electron-phonon coupling and the diagonal elements of the self-energy matrix. To assess the couplings between the phonon modes and the phonon self-energy, we modified the EPW code23 to calculate the off-diagonal elements $\lambda\neq\lambda^{\prime}$ of the phonon self-energy24 and the coupled phonon eigenstates. This mode coupling turns out to be an essential element in strengthening the CDW-PLD formation; see sections Multimode calculations and A Model of a Kohn ladder.
@@
Figure 2: (color online) Decomposition of self-energy to determine the origin of the $q$-vector close to the instability temperature, $\sigma=0.00565$ Ry. Top row (left to right): (a) Phonon dispersions for various modes, numbered with different colors in the legend. (b) Susceptibility, $\chi(q)$ (dashed red line, right-axis)), and $k$-averaged electron-phonon coupling $G(q)$ (remaining curves, left axis). (c) $G(q)\chi(q)$. (d) Diagonal self-energy $\Pi^{\prime}_{\lambda\lambda}$. (e,f) Real (e) and imaginary (f) parts of the off-diagonal self-energy $\Pi^{\prime}_{3\lambda}$, where 3 denotes the soft-mode phonon mode (green). (g) $1/\Omega$. and (h) $\Pi$. Positions of the soft-mode $\bf Q_{CDW}$-vector (solid black vertical lines) and $\bf q=2/3\,\overrightarrow{\Gamma M}$ (dashed black vertical lines) are marked.
@@
Figure 1 summarizes the evolution of the phonon spectra (a-c) and lattice parameters (d) for various $\sigma$ values. Strong $T$-dependence of the phonon frequencies and the appearance of phonon softening are clear evidence that the interactions are inseparably incorporated in the DFT calculations. Strong $T$-dependence of the atomic positions and lattice constants in the numerical results shown in Figure 1(d) suggest the importance of multiphonon processes (noninteracting phonons do not cause thermal lattice expansion). The phonon self-energy here is calculated non-perturbatively in terms of dressed phonons, which leaves open the questions: How does one extract the bare phonon frequencies? And is the DFT based self-energy consistent with the field theory renormalization of the bare phonon frequencies?
Figure 2 presents the phonon spectrum along the $\Gamma-M$ line in panel (a), while the remaining frames provide results for various components of the calculations to gain insight into the nature and origin of the CDW instability. Modes 3, 6, and 8 have predominantly longitudinal symmetry and interact strongly with each other. Modes 2 and 7 are predominantly transverse, but since the modes are strictly longitudinal or transverse only along the high-symmetry lines such as $\Gamma-M$, these longitudinal and transverse modes interact away from the high-symmetry lines. These are the only modes with strong electron-phonon coupling, and therefore, we will we focus mainly on these modes. The sharp dip in the phonon dispersion of mode 3 in panel (a) is the Kohn anomaly that drives the CDW instability50.
@@
To reconstruct the dynamical matrix in the presence of electron-phonon coupling, we start with a general framework where multiple phonon modes interact with multiple electron bands in the vicinity of the Fermi energy. [Mode and band are used here in the Born-Oppenheimer sense where the ionic and electronic degrees of freedom are decoupled.] The electron-phonon interaction is handled perturbatively, see SI for details.
@@
In the multi-mode phonon basis31, the renormalized phonons are given by the poles of the interacting Green’s function:
@@
where $\boldsymbol{\omega}_{\boldsymbol{q}}$ is a diagonal matrix with elements $\omega_{\lambda,\boldsymbol{q}}$ in the phonon eigenmode space describing the bare-phonon frequencies and ${\boldsymbol{\Pi}}(\boldsymbol{q},\omega)$ is the selfconsistent electron-phonon self-energy given by,
@@
$G_{\lambda,\lambda^{\prime}}(n,\boldsymbol{k};m,\boldsymbol{k}+\boldsymbol{q})=\mathfrak{g}_{n,m,\lambda}(\boldsymbol{k},\boldsymbol{k}+\boldsymbol{q})\,\mathfrak{g}_{m,n,\lambda^{\prime}}(\boldsymbol{k}+\boldsymbol{q},\boldsymbol{k})$ (3)
@@
and $g_{n,m,\lambda}(\boldsymbol{k};\boldsymbol{k}+\boldsymbol{q})$ is the microscopic complex electron-phonon coupling given by31
$g_{n,m,\lambda}(\boldsymbol{k};\boldsymbol{k}+\boldsymbol{q})=\langle\Psi_{\boldsymbol{k}+\boldsymbol{q},m}|\nabla_{\boldsymbol{q}}V_{\boldsymbol{q}}.\boldsymbol{\xi}_{\lambda\boldsymbol{q}}|\Psi_{\boldsymbol{k},n}\rangle.$ (5)
where $|\Psi_{\boldsymbol{k},n}\rangle$ are the electron Bloch states, see LABEL:SM2 for details. We also note that $g_{n,m,\lambda}(\boldsymbol{k};\boldsymbol{k}+\boldsymbol{q})=g_{m,n,\lambda}^{*}(\boldsymbol{k}+\boldsymbol{q};\boldsymbol{k})$ by unitarity. Inhomogeneous $\boldsymbol{k}$ dependence in Eq.(5) and hence in Eq.(3) is the result of the anisotropy and the sensitivity of the Bloch wave functions to the crystal structure.
In NbSe2, the relevant sections of the Fermi surface responsible for the strong electron interactions are in the K-pockets near the zone boundaries. In NbSe2, since only one electronic band intersects the Fermi surface, we can neglect the $n,m$ indices. The second term on the right-hand side of Eq.(2) is closely related to the Lindhard electronic susceptibility (dropping the band index n and considering the spin degeneracy factor 2),
which we calculate using Maximally Localized Wannier Functions (MLWFs) 75. We further simplify the computation by taking the adiabatic limit ($\omega\to 0$). [Non-adiabatic corrections in NbSe2 are found to be small.] Effects of the $\boldsymbol{k}$ dependence of $g_{n,m,\lambda}(\boldsymbol{k};\boldsymbol{k}+\boldsymbol{q})$ can be seen in $\Pi_{\lambda,\lambda^{\prime}}$ evaluated using the limiting cases given by Eqs. (2) and (7).
Note that $\Pi_{\lambda\lambda^{\prime}}$ depends on the bare phonon frequencies, $\omega_{\lambda\boldsymbol{q}}$, but that an exact solution for the dressed phonons can be obtained in principle by replacing self-consistently the bare terms by the dressed terms 38, 60. Below, we will refer to the replacement of $\omega_{\lambda\boldsymbol{q}}$ by $\Omega_{\lambda\boldsymbol{q}}$ as the Hedin model.
In the limit of negligibly weak lattice potential, the Bloch states can be approximated by plane waves, so that $g_{n,m,\lambda}(\boldsymbol{k};\boldsymbol{k}+\boldsymbol{q})$ in Eq.(5) becomes dependent only on the momentum exchanged with the phonons34. The phonon self-energy in Eq.(2) can then be written as:
@@
where $G_{\lambda,\lambda^{\prime}}(\boldsymbol{q})$ matrix elements are the single mode version of Eq.(3) in the plane wave approximation. Eq.(7) also represents the dirty limit phonon self-energy as discussed in Section What determines $\boldsymbol{Q}_{CDW}$?.
We now discuss how the $\boldsymbol{k}$ dependence of $g_{n,m,\lambda}(\boldsymbol{k};\boldsymbol{k}+\boldsymbol{q})$ drives an unconventional phonon softening and CDW formation distinct from the Peierls mechanism35. Figure 2(a) shows the DFPT phonon dispersions6 calculated at a temperature slightly above $T_{CDW}$. The CDW transition is accurately predicted at the wavevector $\boldsymbol{Q}_{CDW}\simeq 2\,{\overrightarrow{\Gamma M}}/3$ which agrees with the corresponding experimental value, see LABEL:SM2 for details. It is convenient to introduce a new quantity here: